OpenStudy (jaweria):

Can someone please help me with Calculus 1 problems ? :)

sam (.sam.):

What's the problem?

OpenStudy (jaweria):

its \[f ' (x) = 2x ^{2/3}+4x ^{5/3}-3x+\sqrt{5}\]

OpenStudy (jaweria):

Find f (x) when f ' (x) is given

sam (.sam.):

You have to do that by integration, because differentiation is f(x)-->f'(x) And so integration will be f'(x)-->f(x) \[\int\limits( 2x ^{2/3}+4x ^{5/3}-3x+\sqrt{5})dx\]

sam (.sam.):

You can apply power rule for each term, can you do that?

OpenStudy (jaweria):

can you please guide me little bit more?

sam (.sam.):

There are 4 terms here, each can be used by power rule separately, by applying the power rule formula which is: \[\large \int\limits x^n~dx=\frac{x^{n+1}}{n+1}\] ------------------------------------------------ You get: \[\int\limits \left(2 x^{2/3}+4 x^{5/3}-3 x+\sqrt{5}\right) \, dx=\frac{6 x^{5/3}}{5}+\frac{3 x^{8/3}}{2}-\frac{3 x^2}{2}+\sqrt{5} x\] Do you agree?

OpenStudy (jaweria):

I m trying to understand that How you got 6x^5/3/5+3x^8/3/2-3x^2/2+radical of 5x

OpenStudy (jaweria):

if you dont mind just to make my life easier, can you please just show me little bit detail work for 6x^5/3/5 then I will try to figure out rest.

sam (.sam.):